A tiger leaps horizontally out of a tree that is 3.40 m high. If he lands 5.20 m from the base of the tree, calculate his initial speed vo. (Neglect any effects due to air resistance.) o2.2 m/s
Answer #1

As tiger leaps horizontally, it's initial velocity in vertical direction is zero. Time for vertical displacement is given by

t =√2h/g ......1

where h is height of tree and g is acceleration due to gravity.

Velocity V in horizontal direction remains constant, hence horizontal displacement

d = V t using value of t from equation 1

V = d/(√2h/g) = 5.2(m)/(√2*3.4(m)/9.81(m/s^2) )

= 6.24 m/s